# A Geometric Unification of Generative AI with Manifold-Probabilistic Projection Models

### v1

@@ 5.1 Results
MNIST results: For the MNIST dataset, we set the latent space dimension to 18 and the additive noise to $\epsilon=0.4$. To calculate FID, we trained an MNIST classifier and computed an embedding distribution for each class. After reconstructing a degraded digit, we classified it and compared its embedding with the corresponding pre-computed class distribution. Table 1 reports the mean SSIM and FID metrics. Our method consistently outperforms both DAE and LDM baselines across all degradation types in terms of FID scores. Notably, DAE occasionally achieved higher SSIM values, although its visual results were inferior. Fig. 6 illustrates restoration examples for Gaussian noise, elastic deformation, and downsampling. Additional experiment included the generation of digits from a pure noise. We generated 200 images from random Gaussian noise and managed to obtain realistic digits (FID=19.5) as can be seen in Fig. 7.
Figure 8: Missing pixels. From left to right: degraded, DAE, Diffusion, LMPPM, original.
Figure 9: Random scribbles. From left to right: degraded, DAE, Diffusion, LMPPM, original.
Figure 10: Gaussian noise, $\sigma=0.3$. From left to right: degraded, DAE, Diffusion, LMPPM, original.
Figure 11: Over sharpening. From left to right: degraded, DAE, Diffusion, LMPPM, original.
Figure 12: Gradual reconstruction of missing pixels degradation.
@@
Noise 0.5 Noise 0.7
SSIM $\uparrow$ FID $\downarrow$ SSIM $\uparrow$ FID $\downarrow$
DAE 0.84 11.18 0.71 67.13
LDM 0.79 13.36 0.67 66.88
LMPPM (ours) 0.72 9.73 0.67 9.58
@@
SCUT-FBP5500 Results: Figs. 8910, and 11 present restoration results in facial images under different types of degradation: random missing pixels, random scribbles, strong Gaussian noise, and over sharpening. The quantitative evaluation in Table 2 supports these visual findings, with our method consistently achieving lower FID scores in all degradation scenarios. For these experiments, we set the latent dimension at 1024 and the additive noise at $\epsilon=0.2$. Although the DAE baseline achieves higher SSIM values in certain cases, the perceptual fidelity of its reconstructions remains inferior to that of our approach. Fig. 12 illustrates the reconstruction after four iterations. Finally, in Fig. 13, although the reconstructed face differs slightly from the original, our method yields a result that exhibits higher perceptual realism and lies closer to the natural face manifold, while alternative methods produce reconstructions with a more artificial appearance.
@@ 6 Summary and Conclusions
In this work, we utilize a (denoising) autoencoder in conjunction with the distance function. Providing an approach where both $F$ and $G$ define the manifold while maintaining their coupling to the distance function $\mathcal{D}$ from it. However, due to potential errors in the outputs of the three networks $G$, $F$ and $\mathcal{D}$, especially when $x$ is far from the manifold, this vector field is not exact. Therefore, rather than applying a single-step (weighted) projection onto the manifold, we proceed iteratively, advancing in small steps along the noisy vector field. We are currently exploring an analogous approach where VAE and GAN are coupled with the distance function. A key practical advantage of our approach is its application in the latent space. This dimensionality reduction significantly enhances the accuracy of the distance function, thereby improving restoration and generation results. Indeed, as shown in our experiments (Section  5.1), comparisons with other leading methods indicate the superior performance of our methods, particularly under severe distortions for different data sets and different distortions.
@@ Acknowledgments
Appendix A Appendix: Detailed Theory
A.1 Kernel Method
Detailed computation of Eq. 2.5
$P_{\text{non-u}}(x)$ $=\int_{\mathcal{M}}P(x,y=G(z))dy=\int_{\mathcal{M}}P(x|y=G(z))P(y)dy$
$=\int_{\mathbb{R}^{d}}P(x|G(z))P(G(z))\sqrt{g}dz=\int_{\mathbb{R}^{d}}P(x|G(z))P(z)dz,$
where the subscript "non-u" stands for non-uniform. Here $dy=\sqrt{g}dz$ is the manifold’s volume element, where $g=\det G$ and $G_{\mu\nu}=\sum_{i=1}^{D}J^{i}_{\mu}J^{i}_{\nu}$ is the induced metric, with the Jacobian of the embedding map given by $J^{i}_{\mu}=\partial G^{i}(z)/\partial z_{\mu}$. In the last equality, we use the identity $P(G(z))=P(z)(\sqrt{g})^{-1}$.
@@ A.2 Score Function
$s_{\text{non-u}}(x)=\nabla_{x}\log{P}_{\text{non-u}}\approx\nabla_{x}\log\hat{P}_{\text{non-u}}=:\hat{s}_{\text{non-u}}(x).$
@@
$\hat{s}_{\text{non-u}}(x)=-\frac{1}{2\sigma_{d}^{2}}\left(x-\bar{G}(x)\right),$ (16)
@@
[…]
Specifically,
$\hat{s}_{\text{non-u}}(x)=\frac{1}{\hat{P}_{\text{non-u}}}\nabla_{x}\hat{P}_{\text{non-u}}=\frac{1}{\hat{P}_{\text{non-u}}}\nabla_{x}\left(\int P(x\mid G(z)){P}_{\text{ker}}(z)dz\right).$
Now,
[…]
The integral of $z$ in the computation of $\bar{G}_{\alpha}(x)$ is approximated by randomly sampling the normal distribution centered around the training point $z_{\alpha}$. Explicitly, we approximate the mean using and average over $n$ samples from $P_{\text{ker}}$
[…]
where $\alpha$ denotes an index in the training set (see Fig. 2). The calculation of $\bar{G}_{\alpha}(x)$ requires evaluating $P_{\text{non-u}}(x)$ in the denominator. In particular, we approximate
[…]
Note that in the computation of $\bar{G}_{\alpha}$ all constant factors $Q_{d},\ Q_{\text{ker}}$ and $\frac{1}{n}$, are canceled between the numerator and the denominator.
Appendix B Appendix: Detailed Experimental Setup
@@ B.2 Degradations
We apply six degradation types to simulate real-world image corruption scenarios. Each degradation is applied at three severity levels (mild, intermediate, and severe) to test the robustness of restoration methods:
Degradation Mild Severe
Gaussian noise ($\sigma$) 0.2 0.3
Elastic deformation ($\sigma$) 1.5 1.1
Super-resolution (downsampling factor) 0.5 0.35
Missing Pixels 0.04 0.1
Number of Scibbles 13 20
Over Sharpening 10 18
@@
  • Missing Pixels: Set black patches with some coverage portion;
  • Scribbles Add $n$ random scribbles with random colors
@@ B.3 Model Architectures
Synthetic data for MPPM experiments use MLP-based networks with a latent dimension of 8, selected based on the low intrinsic dimensionality of these manifolds:
@@
Encoder $3\to 64\to 32\to 16\to 8$ with ReLU
Decoder $8\to 16\to 32\to 64\to 3$ with ReLU
Distance Network $8\to 64\to 32\to 16\to 1$ with ReLU, dropout=0.2
@@
Distance Network $1024\to 100\to 50\to 20\to 1$ with ReLU, dropout=0.2
Table 8: Network architectures for SCUT-FBP5500 experiments. Skip connections connect corresponding Encoder and Decoder layers through concatenation. The encoder and decoder blocks are represented as parameterized functions (shown in italic font), where $C_{in}$, $C_{out}$, and $C_{skip}$ represent the number of input, output, and skip connection channels respectively. Abbreviations: Conv = Conv2d (kernel=3, padding=1), BN = BatchNorm2d, LReLU = LeakyReLU(0.2), MPool = MaxPool2d(2), ConvT = ConvTranspose2d(kernel=2, stride=2), Cat = Concatenation. The bottleneck dimension is 1024.
@@ Appendix C Training and Evaluation
Early stopping Patience: 100 epochs
@@
Learning rates AE: $1\times 10^{-3}$, Distance network: $1\times 10^{-5}$, LDM: $1\times 10^{-3}$
@@
Training epochs MNIST: 100, SCUT-FBP5500: 75
Loss functions DAE: L2, LDM: MSE, LMPPM: Composite loss 14
Early stopping Patience: 8 epochs

### v2

@@ 5.1 Results
[…]
Fig. 8 illustrates restoration examples for Gaussian noise, elastic deformation, and downsampling. Additional experiment included the generation of digits from a pure noise. We generated 200 images from random Gaussian noise and managed to obtain realistic digits (FID=19.5) as can be seen in Fig. 9.
Figure 10: Missing pixels (0.08). From left to right: degraded, DAE, Diffusion, LMPPM, original.
Figure 11: Images with 13 scribbles. From left to right: degraded, DAE, Diffusion, LMPPM, original.
Figure 12: Noise $\sigma=0.3$. From left to right: degraded, DAE, Diffusion, LMPPM, original.
Figure 13: Over sharpening (18). From left to right: degraded, DAE, LDM, LMPPM, original.
Figure 14: Gradual reconstruction of missing pixels (0.08) degradation.
@@
SCUT-FBP5500 Results: Figs. 101112, and 13 present restoration results in facial images under different types of degradation: random missing pixels, random scribbles, strong Gaussian noise, and over sharpening. The quantitative evaluation in Table 2 supports these visual findings, with our method consistently achieving lower FID scores in all degradation scenarios. For these experiments, we set the latent dimension at 1024 and the additive noise at $\epsilon=0.2$. Although the DAE baseline achieves higher SSIM values in certain cases, the perceptual fidelity of its reconstructions remains inferior to that of our approach. Fig. 14 illustrates the reconstruction after four iterations. Finally, in Fig. 15, although the reconstructed face differs slightly from the original, our method yields a result that exhibits higher perceptual realism and lies closer to the natural face manifold, while alternative methods produce reconstructions with a more artificial appearance.
CelebA Dataset result We evaluate our method on the CelebA-HQ-256 dataset, adopting the same architecture used for SCUT-FBP5500. The qualitative and quantitative results are presented in Figs. 16 and 17, and Table 3. Our approach achieves a markedly lower FID score, while maintaining SSIM values comparable to the competing methods.
We further compared our method to DiffBIR Lin et al. (2024). DiffBIR tackles blind image restoration using two stages: degradation removal and information regeneration. The first stage removes degradations and produces a high-fidelity but often over-smoothed intermediate result, while the second stage regenerates realistic textures and details. For completeness, we conducted three experiments: (i) DiffBIR after its first stage only, (ii) full DiffBIR, and (iii) our LMPPM followed by DiffBIR’s second stage. The results are shown in Fig. 18 and in the bottom panel of Table 3.
As can be seen, LMPPM outperforms the first stage of DiffBIR both visually (second and third columns from the left) and quantitatively, especially under the missing-pixel and scribble degradations. The output of DiffBIR’s second stage is realistic and perceptually high-quality. Notably, the full DiffBIR model (fourth column from the left of Fig. 18) performs well in removing Gaussian noise (first row), even though the reconstructed image differs from the original image (right column). The best performance is achieved by applying our LMPPM followed by DiffBIR’s second stage (second column from the right), indicating that our blind degradation-removal module provides a strong foundation for high-quality restoration.
Figure 16: Different degradations applied to the CelebA-HQ-256 dataset. From top to bottom: noise ($\sigma=0.3$), missing pixels (0.1), scribbles (22), over sharpening (12). From left to right: degraded, DAE, LDM, LMPPM (ours), and original.
Figure 17: Excessive Gaussian noise ($\sigma=0.3$) applied on CelebA-HQ-256 dataset. From left to right: degraded, DAE, LDM, LMPPM and original.
Figure 18: Comparison to DiffBIR method Lin et al. (2024). From top to bottom: noise ($\sigma=0.3$), missing pixels (0.1), scribbles (22), over sharpening (12). From left to right: degraded, DiffBIR stage1, LMPPM, DiffBIR (stage1 + stage2), LMPPM+DiffBIR stage2, original
Table 3: Quantitative results on the CelebA-HQ-256 dataset, compared also to the DiffBIR method Lin et al. (2024).
Noise 0.3 Scribbles 22 Miss Pixels 0.1 Sharpen 12
SSIM $\uparrow$ FID $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ SSIM $\uparrow$ FID $\downarrow$
DAE 0.694 43.05 0.817 54.75 0.762 49.23 0.719 46.64
LDM 0.663 34.99 0.793 42.73 0.757 41.38 0.724 34.54
LMPPM (ours) 0.707 23.92 0.757 30.69 0.671 25.63 0.714 28.25
DiffBIR 0.70 24.09 0.69 42.95 0.58 42.52 0.76 31.95
DiffBIR stage1 0.67 28.52 0.71 43.85 0.69 44.01 0.89 28.55
LMPPM+DiffBIR stage2 0.68 22.64 0.69 30.68 0.63 23.21 0.72 25.69
@@ 6 Summary and Conclusions
In this work, we utilize a (denoising) autoencoder in conjunction with the distance function. Providing an approach where both $F$ and $G$ define the manifold while maintaining their coupling to the distance function $\mathcal{D}$ from it. However, due to potential errors in the outputs of the three networks $G$, $F$ and $\mathcal{D}$, especially when $x$ is far from the manifold, this vector field is not exact. Therefore, rather than applying a single-step (weighted) projection onto the manifold, we proceed iteratively, advancing in small steps along the noisy vector field. We are currently exploring an analogous approach where VAE and GAN are coupled with the distance function. A key practical advantage of our approach is that it operates in latent space. This dimensionality reduction substantially improves the accuracy of the learned distance function, which in turn enhances both restoration and generation performance. Indeed, as demonstrated in Section 5.1, comparisons with leading baselines show that our method consistently outperforms competing approaches, particularly under severe distortions across multiple datasets and degradation types.
@@ Acknowledgments
Appendix A Detailed Theory
A.1 Manifold Likelihood Derivation
Here we derive Eq. (5). Let $x$ denote a general point in the ambient space, and let $x^{\prime}$ be a clean data point. We denote the probability density of clean images by $P_{c}(x^{\prime})$. Then,
$P_{{\text{non-u}}}(x)$ $=\int_{\mathcal{M}}P(x,x^{\prime})dx^{\prime}=\int_{\mathbb{R}^{D}}P(x\mid x^{\prime})P_{c}(x^{\prime})dx^{\prime}$
$=\int_{\mathbb{R}^{D}}P(x\mid x^{\prime})\left(\int_{\mathcal{M}}P(G(z))\delta(x^{\prime}-G(z))dV_{\mathcal{M}}\right)dx^{\prime}$
$=\int_{\mathcal{M}}P(x\mid x^{\prime}=G(z))P(G(z))\underbrace{\sqrt{g}dz}_{dV_{\mathcal{M}}}=\int_{\mathbb{R}^{d}}P(x\mid G(z))P(z)dz.$
Here $dV_{\mathcal{M}}=\sqrt{g}dz$ is the manifold’s volume element, where $g=\det G$ and $G_{\mu\nu}=\sum_{i=1}^{D}J^{i}_{\mu}J^{i}_{\nu}$ is the induced metric, with the Jacobian of the embedding map given by $J^{i}_{\mu}=\partial G^{i}(z)/\partial z_{\mu}$. In the last equality, we use the identity $P(G(z))=P(z)(\sqrt{g})^{-1}$.
@@ A.2 Score Function
Recall the score function
$s(x)=\nabla_{x}\log{P_{{\text{non-u}}}}(x)\approx\nabla_{x}\log\hat{P}_{{\text{non-u}}}(x)=:\hat{s}(x).$
@@
$\hat{s}(x)=-\frac{1}{2\sigma_{d}^{2}}\left(x-\bar{G}(x)\right),$ (22)
@@
$\bar{G}_{\alpha}(x)=\frac{1}{\hat{P}_{{\text{non-u}}}(x)Q_{\text{ker}}}\int\left[G(z)P(x\mid G(z))\exp\left({-\frac{\|z-z_{\alpha}\|^{2}}{2\sigma_{\text{ker}}^{2}}}\right)\right]dz.$ (23)
We approximate the integral over $z$ in (23) for $\bar{G}_{\alpha}(x)$ by Monte Carlo sampling from the normal distribution centered at the training point $z_{\alpha}$. Specifically, we estimate the mean by averaging over $n$ samples.
$\int\left[G(z)P(x\mid G(z))\exp\left({-\frac{\|z-z_{\alpha}\|^{2}}{2\sigma_{\text{ker}}^{2}}}\right)\right]dz\approx\frac{1}{Q_{d}}\frac{1}{n}\sum_{z_{i}\in\mathcal{N}(z_{\alpha},\sigma_{\text{ker}}^{2})}G(z_{i})\exp\left({-\frac{\|{x}-G(z_{i})\|^{2}}{2\sigma_{d}^{2}}}\right),$ (24)
where $\alpha$ denotes an index in the training set. The calculation of $\bar{G}_{\alpha}(x)$ requires evaluating $\hat{P}_{\text{non-u}}(x)$ in the denominator. In particular, by (8)
$\hat{P}_{\text{non-u}}=\frac{1}{Q_{\text{ker}}}\int\left[P(x\mid G(z))\exp\left({-\frac{\|z-z_{\alpha}\|^{2}}{2\sigma_{\text{ker}}^{2}}}\right)\right]dz\approx\frac{1}{Q_{\text{ker}}Q_{d}}\frac{1}{n}\sum_{z_{i}\in\mathcal{N}(z_{\alpha},\sigma_{\text{ker}}^{2})}\exp\left({-\frac{\|{x}-G(z_{i})\|^{2}}{2\sigma_{d}^{2}}}\right).$ (25)
Substituting (24) and  (25) in (23) yields
$\bar{G}_{\alpha}(x)\approx\frac{\sum_{z_{i}\in\mathcal{N}(z_{\alpha},\sigma_{\text{ker}}^{2})}G(z_{i})\exp\left(-{\|{x}-G(z_{i})\|^{2}/2\sigma_{d}^{2}}\right)}{\sum_{z_{i}\in\mathcal{N}(z_{\alpha},\sigma_{\text{ker}}^{2})}\exp\left({-\|{x}-G(z_{i})\|^{2}/2\sigma_{d}^{2}}\right)}.$ (26)
Note that all constant factors $Q_{d},\ Q_{\text{ker}}$ and $\frac{1}{n}$, are canceled between the numerator and the denominator.
A.3 The Tweedie Formula
The “flow” equations (13) and (15) are the Tweedie formulas for the corresponding probability functions where (13) follows
$P_{d}({x})=\frac{1}{Q_{d}}\exp\left({-\frac{\alpha}{2}{\mathcal{D}^{2}_{\mathcal{M}}({x})}}\right),$ (27)
and (15) follows
$P({x})=\frac{1}{Z}P_{d}({x})\sum_{\alpha\in S}\int_{\mathbb{R}^{d}}\exp\left({-\frac{\beta}{2}{\|{x}-{G}(z)\|^{2}}}\right)\exp\left({-\frac{\|z-z_{\alpha}\|^{2}}{2\sigma_{\text{ker}}^{2}}}\right)dz,$ (28)
where $Z$ is a normalization factor. One can easily verify from (13) that we decrease the distance to the manifold along the flow. Indeed,
$\mathcal{D}_{\mathcal{M}}(x^{n+1})$ $=\mathcal{D}_{\mathcal{M}}(x^{n}-\epsilon\mathcal{D}_{\mathcal{M}}(x^{n})\nabla_{x}\mathcal{D}_{\mathcal{M}}(x^{n}))$
$=\mathcal{D}_{\mathcal{M}}(x^{n})-\epsilon\nabla\mathcal{D}_{\mathcal{M}}(x^{n})\cdot\mathcal{D}_{\mathcal{M}}(x^{n})\nabla_{x}\mathcal{D}_{\mathcal{M}}(x^{n})+O(\epsilon^{2})=(1-\epsilon)\mathcal{D}_{\mathcal{M}}(x^{n})+O(\epsilon^{2}),$
where we used the fact that the distance function is a solution of the Eikonal equation $\|\nabla_{x}\mathcal{D}_{\mathcal{M}}(x)\|^{2}=1$.
Appendix B Detailed Experimental Setup
@@ B.1 Notation and Abbreviations
In the following tables we summarize the abbreviations and datasets through the paper.
@@
CelebA-HQ-256 30,000 high-quality celebrity face images ($256\times 256$)
@@ B.2 Degradations
We apply six degradation types to simulate real-world image corruption scenarios. Each degradation is applied at two severity levels (intermediate, and severe) to test the robustness of restoration methods.
@@
  • Missing Pixels: Set black patches with some coverage portion
  • Scribbles: Add $n$ random scribbles with random colors
@@
Degradation Mild Severe
Gaussian noise ($\sigma$) MNIST 0.7 0.85
Gaussian noise ($\sigma$) Faces 0.2 0.3
Elastic deformation ($\sigma$) 2.3 1.8
Super-resolution (downsampling factor) 0.5 0.35
Missing Pixels 0.04 0.1
Number of Scribbles 6 22
Over Sharpening 8 18
@@ B.3 Model Architectures
Synthetic data for MPPM experiments use MLP-based networks with a latent dimension of 1, selected based on the low intrinsic dimensionality of this manifold:
@@
Encoder $3\to 64\to 64\to 1$ with ReLU
Decoder $1\to 64\to 64\to 3$ with ReLU
Distance Network $3\to 16\to 8\to 4\to 1$ with ReLU
@@
Table 9: Network architectures for SCUT-FBP5500 and CelebA-HQ-256 experiments. Skip connections connect corresponding Encoder and Decoder layers through concatenation. The encoder and decoder blocks are represented as parameterized functions (shown in italic font), where $C_{in}$, $C_{out}$, and $C_{skip}$ represent the number of input, output, and skip connection channels respectively. Abbreviations: Conv = Conv2d (kernel=3, padding=1), BN = BatchNorm2d, LReLU = LeakyReLU(0.2), MPool = MaxPool2d(2), ConvT = ConvTranspose2d(kernel=2, stride=2), Cat = Concatenation. The bottleneck dimension is 1024.
@@
Distance Network $1024\to 300\to 100\to 30\to 1$ with ReLU, dropout=0.2
@@ Appendix C Training and Evaluation
Early stopping Patience: 100
@@
Learning rates AE: $1\times 10^{-3}$, Distance network: $1\times 10^{-3}$ (MNIST $1\times 10^{-5})$, LDM: $1\times 10^{-3}$
@@
Training epochs MNIST: 200, SCUT-FBP5500: 75
Loss functions DAE: L2, LDM: MSE, LMPPM: Composite loss 20
Early stopping Patience: 10 epochs